OpenStudy (anonymous):

okay im stuck here -(3x - 2) = -4(x + 1) - 4 -3x+2=-4x-3-4 do i combine those negatives on the same side ?

OpenStudy (austinl):

You want all terms with variables on one side, and constants on the other.

OpenStudy (anonymous):

i know but you see where ei have the -3 and -4 on one side? I'm not sure do i combine those two before i move on to the next step

OpenStudy (austinl):

Yep, -3-4=?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

i got 0.6 or 1.6666666666666.......

OpenStudy (anonymous):

R.H.S=-4x-4-4

OpenStudy (anonymous):

-3X+2=-4X-4-4 -3X+4X=-8-2 X=-10

OpenStudy (austinl):

Right right, you improperly distributed.

OpenStudy (austinl):

\(-(3x - 2) = \color{red}{-4}(x + \color{red}{1}) - 4\) \(-3x+2=-4x\color{red}{-4}-4\) \(-3x-4x=-4-4-2\) \(\color{blue}{x=-10}\)